CC12:

What is the equation of a line that passes through the point (1, 8) and is perpendicular to the line whose equation is y=x2+3 ? MY ANSWER y=-2x+10

CC12:

@Shadow can you check my answer please

silvernight269:

@umm

Shadow:

what is x2

CC12:

slope?

Shadow:

Yeah but that looks like x^2 which is a parabola

Shadow:

that doesnt have a slope, so I assume you wrote the slope incorrectly.

CC12:

i did oops its x/2

CC12:

yes

CC12:

Thanks Shadow the answer was y=-2x=10

Shadow:

Point-Slope Form \[y - y_{1} = m(x - x_{1})\] \[y - 8 = -2(x - 1)\] \[y - 8 = -2x + 2\] \[y = -2x + 10\] Yeah you're correct.

CC12:

Thank you i made an 80 on the quiz cuz i got that question wrong